1. A 0.75-m^3 rigid tank initially contains air whose density is 1.18 kg/m^3. The tank is connected to a high- pressure supply line through a valve. The valve is opened, and air is allowed to enter the tank until the density in the tank rises to 4.95 kg/m^3. Determine the mass of air that has entered the tank.


2. A desktop computer is to be cooled by a fan whose flow rate is 0.40 m^3/min. Determine the mass flow rate of air through the fan at an elevation of 3400 m where the air density is 0.7 kg/m^3. Also, if the average velocity of air is not to exceed 110 m/min, determine the minimum diameter of the casing of the fan.

1
Expert's answer
2021-11-13T02:10:09-0500


1.

Velocity of the ball at maximum height is zero, v=0

"v \n^2\n \u2212u \n^2\n =2aH"

"0\u2212(49) \n^2\n =2\u00d7(\u22129.8)\u00d7H"

H=122.5 m   


2.

v=u+at

0=49−9.8t

t=5s

Total time taken by ball to return to the surface, T=2t=10 s